Archimedes Levers is financed by a mixture of debt and equity. Complete the following: (Do not round intermediate calculations. Enter your rE and rA answers as a percent rounded to 2 decimal places. Round your beta answers to 2 decimal places.)

rE = 20.20 % rD = 7 % rA = 14.92 %
βE = .90 βD = ? βA = ?
rf = 4 % rm = 22 % D/V = .40
Answer #1

As per CAPM,

rD=rf+ βD*(rm-rf)
7=4+βD(22-4)
βD=1/6=0.17

Weight of debt be x

x*7+(1-x)*20.20=14.92

7x+20.20-20.20x=14.92

-13.20x=5.28
x=0.40

Hence debt =40% equity =60%

(0.40*βD)+(0.60βE)=βA

(0.40*1/6)+(0.60*0.90)=0.61